# Unfitted finite element interpolated neural networks

### v1

@@ 4.1.5. Convergence test for a nonlinear equation
Similar to Sect. 4.1.3, we conduct $h$-refinement tests for the nonlinear equation on the flower domain. We use $k_{U}\in\{2,3\}$ and change the background mesh size for each $k_{U}$. The results are shown in Fig. 8. For this more challenging nonlinear problem, the FEINN errors still exhibit the expected convergence rates, aligning with the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines. The NN errors constantly fall below the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines, suggesting a similar accuracy improvement over FEINNs as seen in the Poisson equation experiments.
Figure 8. Convergence of the $L^{2}$ and $H^{1}$ errors with respect to the trial space mesh size for a 2D nonlinear problem. The figure shows results for two orders: $k_{U}=2$ and $k_{U}=3$, distinguished by colours.
@@ 4.2.1. Convergence tests on analytic geometries
Fig. 10 illustrates the decay of $L^{2}$ and $H^{1}$ errors for FEINNs and NNs as the mesh size changes, with different colours representing the two geometries. For reference, the $\left\lVert u-\pi_{h}(u)\right\rVert$ dashed lines, showing the expected slopes, are also included. The FEINN errors clearly achieve the expected convergence rates as they all fall on the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines. The NN errors are well below the dashed lines, suggesting superior accuracy of the NNs, even with complex 3D geometries defined by level-set functions.
@@ 4.2.2. Training on the octopus geometry defined by STL meshes
We present the NN error in Fig. 11(b) and its gradient error magnitude in Fig. 11(c). The NN solution is highly accurate, with point-wise errors below $3\times 10^{-5}$ and gradient error magnitudes below $6\times 10^{-5}$. For comparison, Tab. 1 lists the errors for the FE interpolation of the analytic solution $u$, FEINN and NN solutions. As expected, the FEINN errors are comparable to the $\pi_{h}(u)$ errors, while the NN errors are noticeably smaller than those of $\left\lVert u-\pi_{h}(u)\right\rVert$. Tab. 1 confirms that unfitted FEINNs are able to accurately approximate the Poisson problem on complex 3D geometries. It also demonstrates that, for smooth solutions, the trained NN can achieve higher accuracy than its interpolation, even with highly intricate 3D shapes defined by STL meshes. Additionally, it also opens the possibility of using FEINNs for solving problems in real-world CAE applications.
FE interpolation of $u$ FEINN NN
@@
Table 1. Errors of the FE interpolation of the true solution, FEINN and NN solutions for the 3D Poisson problem on the octopus geometry.
@@ 4.3. Inverse nonlinear problems
We adopt the NN architectures from [Badia2024, Sect. 6.2.1]: both NNs $u_{\mathcal{N}}$ and $\sigma_{\mathcal{N}}$ have two hidden layers with 20 neurons each; the output layer of $\sigma_{\mathcal{N}}$ is equipped with a rectification function $r(x)=|x|+0.01$ to ensure positivity. All the FE spaces, i.e., $u_{\mathcal{N}}$ interpolation or trial space, $\sigma_{\mathcal{N}}$ interpolation space, and test space, are continuous Galerkin spaces of first order on a background mesh consisting of $50\times 50$ quadrilaterals. These FE spaces are standard unfitted spaces without any stabilisation. Similar to forward problems, essential boundary conditions are weakly imposed via Nitsche’s method with $\gamma=0.01$. The training starts with 400 iterations for $u_{\mathcal{N}}$ in step 1, followed by 100 iterations for $\sigma_{\mathcal{N}}$ in step 2. Step 3 compromises three substeps, with 500 iterations each for the coupled training of $u_{\mathcal{N}}$ and $\sigma_{\mathcal{N}}$. The penalty coefficients $\alpha$ are set to 0.01, 0.03 and 0.09 for these substeps, respectively.
@@
The relative errors during training for the state and coefficient solutions corresponding to the results in Fig. 12 are displayed in Fig. 13. The definition of the relative error is the norm of the difference between the true and identified solutions, divided by the norm of the true solution. It is evident from Fig. 13 that the three-step training strategy effectively addresses the inverse nonlinear problem with partial observations. The state solution errors decrease rapidly during step 1 (iterations 1-400), while the coefficient solution errors decrease quickly during step 2 (iterations 401-500). The coupled training in step 3 further reduces the errors in both state and coefficient solutions. For the state, the interpolated NN solution $\pi_{h}(u_{\mathcal{N}})$ stops improving after 700 iterations due to the low-order interpolation of $u_{\mathcal{N}}$. Despite this, $\pi_{h}(u_{\mathcal{N}})$ already provides satisfactory solution, with relative $L^{2}$ and $H^{1}$ errors below $0.1\%$ and $3\%$, respectively. Moreover, the state NN solution $u_{\mathcal{N}}$ continues to improve, achieving an order of magnitude gain in accuracy in terms of the relative $H^{1}$ error. For the coefficient, both the NN solution $\sigma_{\mathcal{N}}$ and its interpolation $\pi_{h}(\sigma_{\mathcal{N}})$ exhibit similar convergence curves, with their relative $L^{2}$ errors steadily decreasing to $2\%$.
@@ 5. Conclusions
We conduct extensive numerical experiments to evaluate the unfitted FEINN method on various 2D and 3D complex geometries for different values of numerical parameters, geometries and PDEs. As opposed to unfitted FEM, small cut cells have negligible impact on the accuracy of both trained NNs and their interpolations, even without using stabilisation techniques such as CutFEM or AgFEM to compute the residual (such stabilisations are required, though, for the computation of discrete dual norms, i.e., preconditioned losses). The $h$- and $p$-convergence tests for the Poisson equation manifest that the trained NNs outperform the FE interpolation of the analytic solution on the same trial space, and their interpolations accomplish the expected convergence rates.
Preconditioning based on dual norms in standard test spaces, given proper stabilisation technique such as CutFEM for the Gram matrix, or in aggregated test spaces significantly accelerates NN training convergence for the Poisson equation, particularly when the solution features a sharp gradient in a localised region of the domain. The nonlinear equation experiments verify that unfitted FEINNs are capable of dealing with more challenging PDEs. The experiments on 3D geometries show that unfitted FEINNs accurately solve PDEs on 3D complex geometries defined implicitly via level-set functions or explicitly through STL surface meshes. This makes the method well-suited for real-world CAE scenarios. The inverse problem experiments exhibit that unfitted FEINNs can solve nonlinear inverse problems with high accuracy and reliability on complex geometries.
Future research could explore several promising directions. First, as discussed in [Badia2024adaptive], to exploit the nonlinear approximability properties of NNs, it is essential to adaptively refine the FE spaces. Enhancing unfitted FEINNs with adaptive training techniques [Badia2024adaptive] will enable us to tackle problems with multiscale features or singularities on complex geometries. Another encouraging research avenue is applying unfitted FEINNs to evolutionary PDEs, especially time-dependent problems on moving domains [Badia2023] and fluid-structure interaction problems with evolving geometries [Massing2015]. In time-dependent problems, the NN trained at the previous time step serves as a good initial guess for the current step, which can clearly improve the training efficiency of unfitted FEINNs.

### v2

@@ 4.1.6. Convergence test for a nonlinear equation
Similar to Sect. 4.1.4, we conduct $h$-refinement tests for the nonlinear equation on the flower domain. We use $k_{U}\in\{2,3\}$ and change the background mesh size for each $k_{U}$. The results are shown in Fig. 10. For this more challenging nonlinear problem, the $\pi_{h}(u_{\mathcal{N}})$ errors still exhibit the expected convergence rates, aligning with the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines. The NN errors constantly fall below the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines, suggesting a similar accuracy improvement over $\pi_{h}(u_{\mathcal{N}})$ as seen in the Poisson equation experiments.
Figure 10. Convergence of the $L^{2}$ and $H^{1}$ errors with respect to the trial space mesh size for the 2D nonlinear problem. The figure shows results for two orders: $k_{U}=2$ and $k_{U}=3$, distinguished by colours.
@@ 4.2.1. Convergence tests on analytic geometries
Fig. 12 illustrates the decay of $L^{2}$ and $H^{1}$ errors for $\pi_{h}(u_{\mathcal{N}})$ and $u_{\mathcal{N}}$ as the mesh size changes, with different colours representing the two geometries. For reference, the $\left\lVert u-\pi_{h}(u)\right\rVert$ dashed lines, showing the expected slopes, are also included. The $\pi_{h}(u_{\mathcal{N}})$ errors clearly achieve the expected convergence rates as they all fall on the $\left\lVert u-\pi_{h}(u)\right\rVert$ lines. The NN errors are well below the dashed lines, suggesting superior accuracy of the NNs, even with complex 3D geometries defined by level-set functions.
@@ 4.2.2. Training on the octopus geometry defined by STL meshes
We present the NN error in Fig. 13(b) and its gradient error magnitude in Fig. 13(c). The NN solution is highly accurate, with point-wise errors below $3\times 10^{-5}$ and gradient error magnitudes below $6\times 10^{-5}$. For comparison, Tab. 1 lists the errors for the FE interpolation of the analytic solution $u$, $\pi_{h}(u_{\mathcal{N}})$ and $u_{\mathcal{N}}$ solutions. As expected, the $\pi_{h}(u_{\mathcal{N}})$ errors are comparable to the $\pi_{h}(u)$ errors, while the NN errors are notably smaller than those of $\left\lVert u-\pi_{h}(u)\right\rVert$. Tab. 1 confirms that unfitted FEINNs are able to accurately approximate the Poisson problem on complex 3D geometries. It also demonstrates that, for smooth solutions, the trained NN can achieve higher accuracy than its interpolation, even with highly intricate 3D shapes defined by STL meshes. Additionally, it also opens the possibility of using FEINNs for solving problems in real-world CAE applications.
$\pi_{h}(u)$ $\pi_{h}(u_{\mathcal{N}})$ $u_{\mathcal{N}}$
@@
Table 1. Errors of the FE interpolation of the true solution, interpolation of the trained NN and NN itself for the 3D Poisson problem on the octopus geometry.
@@ 4.3. Inverse nonlinear problems
We adopt the NN architectures from [Badia2024, Sect. 6.2.1]: both NNs $u_{\mathcal{N}}$ and $\sigma_{\mathcal{N}}$ have two hidden layers with 20 neurons each; the output layer of $\sigma_{\mathcal{N}}$ is equipped with a rectification function $r(x)=|x|+0.01$ to ensure positivity. All the FE spaces, i.e., $u_{\mathcal{N}}$ interpolation or trial space, $\sigma_{\mathcal{N}}$ interpolation space, and test space, are continuous Galerkin spaces of first order on a background mesh consisting of $50\times 50$ quadrilaterals. These FE spaces are standard unfitted spaces without any stabilisation. Similar to forward problems, essential boundary conditions are weakly imposed via Nitsche’s method with $\gamma=10^{-2}$. The training starts with 400 iterations for $u_{\mathcal{N}}$ in step 1, followed by 100 iterations for $\sigma_{\mathcal{N}}$ in step 2. Step 3 compromises three substeps, with 500 iterations each for the coupled training of $u_{\mathcal{N}}$ and $\sigma_{\mathcal{N}}$. The penalty coefficients $\alpha$ are set to 0.01, 0.03 and 0.09 for these substeps, respectively.
@@
The relative errors during training for the state and coefficient solutions corresponding to the results in Fig. 14 are displayed in Fig. 15. The label “NN interp” refers to the interpolated NN solutions, while “NN” refers to the NN solutions themselves. The definition of the relative error is the norm of the difference between the true and identified solutions, divided by the norm of the true solution. It is evident from Fig. 15 that the three-step training strategy effectively addresses the inverse nonlinear problem with partial observations. The state solution errors decrease rapidly during step 1 (iterations 1-400), while the coefficient solution errors decrease quickly during step 2 (iterations 401-500). The coupled training in step 3 further reduces the errors in both state and coefficient solutions. For the state, the interpolated NN solution $\pi_{h}(u_{\mathcal{N}})$ stops improving after 700 iterations due to the low-order interpolation of $u_{\mathcal{N}}$. Despite this, $\pi_{h}(u_{\mathcal{N}})$ already provides satisfactory solution, with relative $L^{2}$ and $H^{1}$ errors below $0.1\%$ and $3\%$, respectively. Moreover, the state NN solution $u_{\mathcal{N}}$ continues to improve, achieving an order of magnitude gain in accuracy in terms of the relative $H^{1}$ error. For the coefficient, both the NN solution $\sigma_{\mathcal{N}}$ and its interpolation $\pi_{h}(\sigma_{\mathcal{N}})$ exhibit similar convergence curves, with their relative $L^{2}$ errors steadily decreasing to $2\%$.
@@ 5. Conclusions
We conduct extensive numerical experiments to evaluate the unfitted FEINN method on various 2D and 3D complex geometries for different values of numerical parameters, geometries and PDEs. As opposed to unfitted FEM, small cut cells have negligible impact on the accuracy of both trained NNs and their interpolations, even without using stabilisation techniques such as CutFEM or AgFEM to compute the residual (such stabilisations are required, though, for the computation of discrete dual norms). The $h$- and $p$-convergence tests for the Poisson equation manifest that the trained NNs outperform the FE interpolation of the analytic solution on the same trial space, and their interpolations accomplish the expected convergence rates. Besides, unfitted FEINNs demonstrate robustness to the choice of Nitsche coefficient, which is a another advantage over unfitted FEMs.
In experiments comparing unfitted FEINNs and PINNs, FEINNs achieve comparable or better accuracy, demonstrate higher robustness to hyperparameter choices, and require significantly less training time on both GPUs and CPUs. We observe that the use of a discrete dual norm of the residual in the loss function, combined with a robust unfitted discretisation, significantly accelerates NN training convergence compared to the Euclidean norm variant and PINNs. This effect is particularly notable for Poisson problems featuring a sharp gradient in a localised region of the domain, substantially reducing computational cost. The nonlinear equation experiments verify that unfitted FEINNs are capable of dealing with more challenging PDEs. The experiments on 3D geometries show that unfitted FEINNs accurately solve PDEs on 3D complex geometries defined implicitly via level-set functions or explicitly through STL surface meshes. This makes the method well-suited for real-world CAE scenarios.
The inverse problem experiments exhibit that unfitted FEINNs can be easily and effectively extended to solve nonlinear inverse problems, delivering highly accurate and robust solutions on complex geometries. This overcomes key limitations of traditional FEMs for inverse settings, which require the use of adjoint solvers, nested loops, and regularisation.
Future research could explore several promising directions. First, as discussed in [Badia2025adaptive], to exploit the nonlinear approximability properties of NNs, it is essential to adaptively refine the FE spaces. Enhancing unfitted FEINNs with adaptive training techniques [Badia2025adaptive] will enable us to tackle problems with multiscale features or singularities on complex geometries. Another encouraging research avenue is applying unfitted FEINNs to evolutionary PDEs, especially time-dependent problems on moving domains [Badia2023] and fluid-structure interaction problems with evolving geometries [Massing2015]. In time-dependent problems, the NN trained at the previous time step serves as a good initial guess for the current step, which can clearly improve the training efficiency of unfitted FEINNs.
@@ Acknowledgments
Declaration of generative AI and AI-assisted technologies in the writing process
During the preparation of this work the authors used ChatGPT in order to improve language and readability. After using this tool/service, the authors reviewed and edited the content as needed and take full responsibility for the content of the publication.